HETAV PATEL

Grade 11,

how to name the compounds of the type like... [Pt(NH 3 )Cl 2 ][PtCl 4 ]??????? please state the procedure??????

how to name the compounds of the type like...


[Pt(NH3)Cl2][PtCl4]???????


please state the procedure??????

Grade:12

1 Answers

ruchi yadav
10 years ago
Pt(NH3)4Cl2]: Pt(IV) -2(two Cl^-) = 2+ hence [Pt(NH3)4Cl2]^2+
[PtCl4]: Pt(II) - 4(four Cl^-) = 2-
It is understood by TM chemists (but not perhaps by beginners!) that once you specify the oxidation state of the central metal atom (from experiment) it is not necessary to specify the charge on the cation or anion.

Thank You
Ruchi
Askiitians Faculty